# Non-stationary Spatio-Temporal Modeling Using the Stochastic Advection-Diffusion Equation

### v1

@@ 4.4 Results
The results are shown in Figure 4. Here, the $x$-axis shows the difference between the timestep when data is observed and the time step when predictions are made $\Delta k_{p}=k_{p}-k_{o}$, where $k_{o}$ is the observed timestep and $k_{p}$ is the prediction timestep. Observations at $k_{o}$ are collected at all locations except the masked area, and predictions are made for all locations at $k_{p}$, except if $k_{p}=k_{o}$, then predictions are made only for the masked area. In Figure 4, the reconstruction error is shown for $\Delta k_{p}=0$ and the forecasting error is for $\Delta k_{p}>0$. For instance, for our timeseries which consist of $12$ timesteps, consider we observe all locations except the masked area at the third timestep ($k_{2}$). Then, the reconstruction is made for the masked area at $k_{2}$ and predictions are made for all locations in the nine future timesteps $0<\Delta k_{p}\leq 9$. In Figure 4, this is added to the evaluation metrics for prediction timesteps $0\leq\Delta k_{p}\leq 9$. If $k_{o}=0$ is observed then $k_{p}=[0,11]$, and if $k_{o}=11$ is observed then only the reconstruction error for $k_{p}=11$ is calculated and errors added to prediction timestep $\Delta k_{p}=0$. This is done for each of the models and an average is taken over the respective $\Delta k_{p}$ made from all combinations of $k_{o}$ and $k_{p}$ and the $20$ realizations in the test data.
@@
Figure 4: Root mean square error (left) and continuous ranked probability score (right) for the reconstruction ($\Delta k_{p}=0$) and forecasting ($\Delta k_{p}>0$) of the different models.
@@ 5.1 Motivation and goal
The goal is to compare the proposed complex non-separable model to a separable model to evaluate how much is gained by a non-separable model. In this application, the true physical process (and the numerical model) clearly has non-separable phenomena such as diffusion and advection, which gives a good platform to evaluate the proposed approach. Previous works have in general employed purely spatial GRFs onboard the AUVs for real-time prediction, interpolation, and monitoring (Fossum et al.,, 2021; Ge et al.,, 2023; Berild et al.,, 2024). However, a strong disadvantage is that the ocean is dynamic, and assuming that the field is constant in time is not realistic for the general duration of an AUV mission, but it does speed up the onboard computations. Furthermore, many of these approaches have used adaptive sampling strategies to choose locations for the AUV to collect measurements, but since our main focus is prediction accuracy, we will generate random paths for the AUV to follow.
@@ 5.2 Estimating the surrogate models
The area of interest spans $1568\,\mathrm{m}\times 1408\,\mathrm{m}$, and we use a discretization of $M=50$ cells in $x$-direction and $N=45$ cells in $y$-direction, which gives a total of $K=M\cdot N=2250$ spatial locations. In addition, a buffer around the field is added to mitigate the effect of the boundaries of the field, and the total area is $M=60$ cells in $x$-direction and $N=55$ cells in $y$-direction. There is also added an advection tapering within this buffer to mitigate the effects of the boundaries even further. This area is large enough to capture the dynamics of the river outlet and the surrounding area, and with the number of timesteps would be unpractical to handle with classical spatio-temporal models. We use May 27, 2021, for estimating the model, termed the training data. This is 144 timesteps, where each time step is 10 min, and we split the dataset into $9$ non-overlapping time segments of $T=9$ timesteps (90 minutes) with one hour between each segment to avoid temporal correlation between them.
We collect the training data from each time segment as $\bm{y}_{1},\ldots,\bm{y}_{m},\ldots,\bm{y}_{9}$, where $\bm{y}_{m}=(y_{m,1,1},\ldots,y_{m,k,t},\ldots,y_{m,K,T})^{\top}$ is the salinity field of $m$th time segment. Since this data has a trend that varies through time and the time segments are assumed to be independent of each other, we remove this trend by subtracting the mean of each time segment from the data as $\tilde{\bm{y}}_{m}=\bm{y}_{m}-\bar{\bm{y}}_{m}$, where $\bar{\bm{y}}_{m}$ is the mean over the time segment for each location in space
$\bar{\bm{y}}_{m}=\frac{1}{T}\sum_{t=1}^{T}\bm{y}_{m,t}.$
Here, $\bm{y}_{m,t}=(y_{m,t,1},\ldots,y_{m,t,K})^{\top}$ is the spatial salinity field at time $t$ for the $m$th time segment. The models are then estimated on the resulting residuals of the detrended SINMOD data with observation model
$\tilde{\bm{y}}_{m}=\mathbf{E}{\bm{w}}_{m}+\bm{\epsilon}_{m},\quad m=1,\ldots,9,$
@@
We consider the proposed flexible stochastic diffusion-advection-reaction model termed NStat-AD and the complex non-stationary separable model termed NStat-Sep described in Section 4.3. In terms of the initial values of the parameters, it is quite easy to make reasonable guesses for the advection parameters as they most likely will be similar to the currents which are generally pointed to the North (given by SINMOD output). The diffusion parameters are harder to suggest but are initially set perpendicular to the advection vector field.
@@ 5.3 Evaluating the surrogate models
The test dataset is comprised of simulation data from SINMOD spanning nine different days: May 28, May 29, May 4, May 10, May 11 in 2021, and June 21, June 22, September 8, and September 9 in 2022. The setup for the test data mirrors that of the training data and includes 72 time segments, each treated as independent from one another and from the training set. However, the test data is not detrended as the training data, as the emulator is expected to predict the salinity field directly from measurements. Therefore, the observation model for the test data is given by
$\bm{y}_{m}=\bm{\mu}+\mathbf{E}{\bm{w}}_{m}+\bm{\epsilon}_{m},\quad m=1,\ldots,72,\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\sigma_{n}^{2}\mathbf{I}),$
@@
where $\bm{\mu}=(\bm{\mu}_{\mathrm{S}},\ldots,\bm{\mu}_{\mathrm{S}})^{\top}$ stacked $T$ times with $\bm{\mu}_{\mathrm{S}}$ which is the spatially varying empirical mean of the training data across all time segments. In real-world applications, it might not be feasible to develop surrogate models that are specifically tailored for the exact dates and durations of upcoming missions. Therefore, testing the models with SINMOD data from different days in this simulated environment serves as an effective way to assess the models’ adaptability and predictive accuracy.
@@
We assume that the AUV moves at a constant speed of $0.5\,\mathrm{m/s}$, with each mission lasting 90 minutes, matching the time segments of the model. The AUV’s trajectories are designed as random walks that minimize path intersections, favor straight-line movement, and remain within the designated area of interest. An example of such a path is depicted in Figure 7(a). In practical scenarios, AUVs would collect data at specific intervals; however, for this simulation, we assume data collection occurs whenever the AUV enters a new grid cell or a new time step begins. This data collection procedure is very similar to the result of data assimilation done with finite volume solution of the SPDE model (Berild et al.,, 2024). For this study, we generated 200 such random paths, and the specific paths employed are illustrated in Figure 7(b).
For each of the 72 time segments and 200 simulated paths, measurements are collected with a measurement error variance of $\sigma_{n}^{2}=0.01$. This measurement error is introduced to compensate for the deterministic and overly smooth outputs of the SINMOD model, simulating the noise typically found in real-world data. In total, this is 14400 different missions, and for each of them the estimated surrogate models are updated with the respective measurements to obtain the posterior distributions. The posterior mean serves as the spatial and temporal point prediction, with accuracy assessed via the RMSE on unobserved points for each mission. We then calculate both the mean and standard deviation of the RMSE across all paths for each dataset. To evaluate the predictive distribution further, the CRPS is computed for each mission. Similarly, the mean and standard deviation of the CRPS are calculated across the 200 paths for each of the 72 datasets.
@@ 6 Discussion
Stationary models are contained in the new model class by using one parameter per coefficient instead of basis functions. The estimated models are interpretable through an initial distribution, and the diffusion tensor and advection vector field. Further, in the advection-dominated simulation study, we do not observe the instability seen by Clarotto et al., (2023) for their FEM approximation. This suggests that FVM is a promising alternative to FEM for this problem.
@@
Optimizing our model presents several challenges and opportunities for improvement. One notable limitation is that the vector field describing the anisotropic diffusion exhibits two equivalent optima, complicating optimization efforts, particularly since this vector field varies spatially. Exploring alternative parameterizations of the vector field that circumvent this issue could be beneficial. In our efforts to optimize the model parameters, we have utilized the ADAM optimizer (Kingma and Ba,, 2017), chosen for its effectiveness in handling vanishing gradients—a common issue in machine learning scenarios where certain parameter gradients are extremely small. Despite its advantages, ADAM is not without flaws, and there may be potential for better performance with other optimizers. We have tested additional optimizers, including Rmsprop, Adagrad, and Adadelta (Ruder,, 2017), but none outperformed ADAM in our studies. We also experimented with AdaHessian (Yao et al.,, 2021), which scales gradients using the diagonal of the Hessian matrix. However, this approach did not enhance the optimization process, likely because the parameters are too interdependent to benefit from scaling by only the diagonal elements. Calculating the full Hessian matrix, while potentially more effective, is prohibitively expensive in computational terms for this model. Exploring these optimization challenges further could yield significant advancements in the performance and applicability of our approach.
@@
Appendix A. Discretizing the advection-diffusion SPDE
@@ A.1 Problem formulation
$\left(\mathbf{H}(\bm{s})\nabla u(\bm{s},t)+\bm{\omega}(\bm{s})u(\bm{s},t)\right)\cdot{\bm{n}}(\bm{s})=0,\quad\bm{s}\in\partial\mathcal{D},\quad t\in\mathcal{T},$
@@ A.2 Temporal discretization
$\phantom{=}+\Delta t\left\{-\left[\kappa(\cdot)^{2}-\nabla\cdot\mathbf{H}(\cdot)\nabla\right]u^{n+1}(\cdot)+\nabla\cdot(\bm{\omega}(\cdot)u^{n+1}(\cdot))\right\}+$

### v2

@@ 4.4 Results
The results are shown in Figure 4. Here, the $x$-axis represents the lag, $\Delta k_{\mathrm{p}}=t_{\mathrm{p}}-t_{\mathrm{o}}$, between the time step when predictions are made, $t_{\mathrm{p}}$, and the time step when data is observed, $t_{\mathrm{o}}$. Reconstructions, $\Delta k_{\mathrm{p}}=0$, only computes RMSE and CRPS for the test area, while forecasts $\Delta k_{\mathrm{p}}>0$ computes RMSE and CRPS for both test and training locations. For each candidate model and for each $\Delta k_{\mathrm{p}}$, we take an average and an empirical standard deviation for the scores corresponding to legal time step combinations and realizations as discussed in Section 4.2.
@@
Figure 4: RMSE (left) and CRPS (right) for the reconstruction ($\Delta k_{p}=0$) and forecasting ($\Delta k_{p}>0$) of the different models. The solid line shows the average and shaded region shows $\pm$ one standard deviation. See Section 4.2 for details.
@@ 5.1 Motivation and goal
The goal is to compare the proposed complex non-separable model to a separable model to evaluate how much is gained by a non-separable model. In this application, the true physical process (and the numerical model) clearly has diffusion and advection, which gives a good platform to evaluate the proposed approach. Previous works have in general employed purely spatial GRFs onboard the AUVs for real-time prediction, interpolation, and monitoring (Fossum et al.,, 2021; Ge et al.,, 2023; Berild et al.,, 2024). However, a strong disadvantage of previous work is that the ocean is dynamic, and assuming that the field is constant in time is not realistic for the typical duration of an AUV mission. Purely spatial models are usually used because they speed up onboard computations compared to spatio-temporal models.
We use paths that are pregenerated randomly, and treat the path as deterministic. The paths could also be chosen adaptively (Berild et al.,, 2024), but since our main focus is to evaluate prediction accuracy, we do not bring in further complications that can make it harder to interpret differences between models.
@@ 5.2 Estimating the surrogate models
The area of interest spans $1568\,\mathrm{m}\times 1408\,\mathrm{m}$, and we use a discretization of $M=50$ cells in $x$-direction and $N=45$ cells in $y$-direction, which gives a total of $K=M\cdot N=2250$ spatial locations. In addition, a buffer around the field is added to mitigate the effect of the boundaries on the marginal variances of the field, and the total size is $M=60$ cells in $x$-direction and $N=55$ cells in $y$-direction. This corresponds to a $160\,\mathrm{m}$ buffer. We also add advection tapering within this buffer to mitigate the effects of the boundaries even further. This area is large enough to capture the dynamics of the river outlet and the surrounding area, and with a number of timesteps that would be impractical to handle with classical spatio-temporal models.
We use May 27, 2021 for estimating the model. We extract the data from SINMOD and add i.i.d. Gaussian noise with variance $\sigma_{\mathrm{obs}}^{2}=0.01$ to mimic that the truth is less smooth than numerical simulations from SINMOD. We call the resulting data as the training data. The data consists of 144 timesteps, where each time step is 10 min, and we split the dataset into $9$ non-overlapping time segments of $T=9$ timesteps (90 minutes) with one hour between each segment. The length of the time segments was chosen because we aim to estimate a surrogate model that works well for missions of length 1.5 hours. We do not use a moving window approach covering all windows of length 1.5 hours, and instead aim to cover a range of different behaviours without considering all possible 135 moving windows of length 9. Direct computations with 144 time steps would not be possible with Cholesky factorizations. In that case, it would be worthwhile to investigate matrix-free approaches such as discussed in Clarotto et al., (2024), but they did not consider problems as large as 144 time steps in the paper.
The training data is the true salinities $\bm{y}_{m}=(\bm{y}_{m,1}^{\top},\ldots,\bm{y}_{m,T}^{\top})^{\top}$, where $\bm{y}_{m,t}=(y_{m,t,1},\ldots,y_{m,t,K})$ are the true salinities for time step $t$ in time segment $m$, for $t=1,\ldots,T$ and $m=1,\ldots,9$. To reduce effects of temporal evolution at a longer time scale than the envisioned mission lengths of $1.5$ hours, we subtract the temporal mean of each time segment from the data as $\tilde{\bm{y}}_{m}=\bm{y}_{m}-\bar{\bm{y}}_{m}$, where $\bar{\bm{y}}_{m}$ is the temporal mean for each location in space,
$\bar{\bm{y}}_{m}=\frac{1}{T}\sum_{t=1}^{T}\bm{y}_{m,t},\quad m=1,\ldots,9.$
The surrogate models are then estimated on the resulting residuals, $\tilde{\bm{y}}_{1},\ldots,\tilde{\bm{y}}_{9}$, using the model
$\tilde{\bm{y}}_{m}=\mathbf{E}{\bm{u}}_{m}+\bm{\epsilon}_{m},\quad m=1,\ldots,9,$
@@
We consider two of the models described in Section 4.3: 1) the proposed flexible stochastic diffusion-advection-reaction model (NStat-AD), and 2) the complex non-stationary separable model (NStat-Sep). As starting values in the optimization, we choose that advection should be similar to the currents, which are generally pointed to the North (given by SINMOD output). The diffusion parameters are harder to choose, but are initially set perpendicular to the advection vector field.
@@ 5.3 Evaluating the surrogate models
The test dataset is comprised of numerical simulation data from SINMOD spanning nine different days: May 4, May 10, May 11, May 28, and May 29 in 2021, and June 21, June 22, September 8, and September 9 in 2022. The setup for the test data mirrors that of the training data, and for each date, there are 9 time segments for a total of 72 time segments across the nine dates. The two estimated surrogate models in Section 5.2, NStat-Sep and NStat-AD, are used to model the test data.
We consider a generic surrogate model and time segment, and write the estimated surrogate model for the truth as
$\bm{\eta}=\hat{\bm{\mu}}+{\bm{u}}+\bm{\epsilon},$ (21)
@@
where ${\bm{u}}$ has the covariance matrix $\mathbf{Q}^{-1}(\hat{\bm{\theta}})$ estimated in Section 5.2, $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\hat{\sigma_{\mathrm{N}}^{2}}\mathbf{I})$ has the estimated variance $\hat{\sigma_{\mathrm{N}}^{2}}$ from Section 5.2, and $\hat{\bm{\mu}}=(\hat{\bm{\mu}}_{\mathrm{S}}^{\top},\ldots,\hat{\bm{\mu}}_{\mathrm{S}}^{\top})^{\top}$ is a known temporally constant mean with a spatially varying $\hat{\bm{\mu}}_{\mathrm{S}}$ that is repeated $9$ times. The spatially varying mean $\hat{\bm{\mu}}_{\mathrm{S}}$ represents the expected spatial pattern, and is needed because the surrogate models in Section 5.2 are estimated based on residuals from a spatially varying mean. We compute $\hat{\bm{\mu}}_{\mathrm{S}}$ as the temporal average, across the 144 time steps in the training data in Section 5.2, for each spatial location. In real-world applications, it is hard to develop surrogate models that are specifically tailored for the exact dates and duration of upcoming missions. Therefore, we use this setup which evaluates applying the surrogate model to different days than used to estimate the surrogate model. This serves as an effective way to assess the surrogate models’ adaptability and predictive accuracy.
@@
We consider an AUV that moves at a constant speed of $0.5\,\mathrm{m/s}$, and we create preplanned paths for the AUV by simulating random walks that minimize path intersections, favor straight-line movement, and remain within the designated area of interest. An example of such a generated path is depicted in Figure 7(a). In practical scenarios, AUVs would collect data at specific intervals; however, for this simulation, we assume data collection occurs whenever the AUV enters a new grid cell or a new time step begins. This data collection procedure is similar to the result of data assimilation done with FVM approach for the spatial SPDE model in Berild et al., (2024). We generate 200 such paths, and use the same paths for each time segment. The paths are illustrated in Figure 7(b). This gives $72\cdot 200=14400$ combinations of time segments and paths. To simplify the description that follows, we term each such combination as a mission.
For each of the 14400 missions, we assume that values are observed exactly. Typically a new measurement variance would need to be estimated from a trial run in the ocean based on empirically observed variation within grid cells (Berild and Fuglstad,, 2023). However, in this case the resolution matches for the surrogate model and the real-world variables. The data collected in mission $i$ is denoted ${\bm{z}}_{i}$ and is of length $n_{\mathrm{obs},i}$, $i=1,\ldots,14400$.
All missions are evaluated separately, and we describe a generic mission and skip the index $i$ for ease of reading. We link the observation to the latent variables through ${\bm{z}}=\mathbf{E}_{\mathrm{obs}}\bm{\eta}$, where the matrix $\mathbf{E}_{\mathrm{obs}}$ links the $n_{\mathrm{obs}}$ observed locations in space and time to $\bm{\eta}$. The multivariate Gaussian surrogate model in Equation (21) is then updated using the observed data. This gives a new multivariate Gaussian distribution $\bm{\eta}|{\bm{z}}$, which describes posterior beliefs about the true states at all the 9 time steps. The posterior mean, $\mathrm{E}[\bm{\eta}|{\bm{z}}]$ serves as the point predictions for all locations in all time steps, and is assessed by computing the RMSE for the unobserved spatio-temporal locations during the mission. Similarily, we compute the an average CRPS across the unobserved spatio-temporal locations based on the marginal distributions from $\bm{\eta}|{\bm{z}}$.
For each of the 72 time segments, we then compute the mean and the standard deviation of the RMSE values and CRPS values across the 200 missions in that time segment. This is done for each surrogate model.
@@ 5.4 Results
5.5 Sensitivity to model choices
Several key choices were made when construction the model: 1) the spatial resolution, 2) the choice of flexibility in the covariance structure, and 3) the treatement of the boundary. Below we discuss the rationale behind the choices, and test sensitivity to some important choices.
The spatial resolution is chosen to match the spatial resolution ($32\,\mathrm{m}\times 32\,\mathrm{m}$) of the SINMOD data, and is the finest resolution resolvable from the data source. In this work we are not interested in evaluating methods for downscaling to higher resolution, and chose to use the same resolution for the surrogate model. If finer-scale models for the AUVs are desired, future work would be required to determine how to reliably downscale to finer resolutions.
The flexibility with $3\times 3$ basis functions for the coefficients in the SPDEs were chosen to give a moderate amount of flexibility that we expected could be estimated from the data. One might wonder if even higher flexibility would further improve the results compared to the stationary model. We repeated the steps also for a model with $5\times 5$ basis functions (NStat-AD B-spline 25). Table 1 shows that the model still performs better than NStat-Sep, but that performance is worse than the less flexible NStat-AD. From Figure 9, it is clear that the more flexible model is consistently worse. This suggests that there may be a need for penalization of the flexibility to avoid overfitting to the training data. This is also supported by the NStat-AD B-spline 25 model having larger maximum likelihood values than the NStat-AD on the training data.
Figure 9: Sensitivity in Figure 8 to changes to model choices.
Table 1: Scores under changes to model choices.
Model Extension RMSE CRPS ($10^{-1}$)
NStat-Sep - 1.376 9.05
NStat-AD - 0.947 5.06
NStat-AD Buffer 10 Larger buffer zone 0.940 5.12
NStat-AD B-spline 25 More flexibility in covariance 1.026 5.69
Alternative boundary choices such as fixed values on the boundary, and periodic boundary conditions are undesirable for the application. Fixed values on the boundary gives variances close to zero, and periodic boundary conditions allows advection to transport small/large values from one side of the domain to the other. Additionally, it is not straightforward to mitigate such effects in a non-stationary model since there is no natural extension of the model to the entire $\mathbb{R}^{2}$ without knowing the spatially varying coefficients also outside the domain $\mathcal{D}$. In our opinion, the no-flow boundary conditions coupled with the tapering of advection offers most reasonable approach to mitigate undesirable behaviour. To investigate the sensitivity to the choice of using a buffer of 5 cells ($160\,\mathrm{m}$), we repeated the steps for a model (NStat-AD Buffer 10) with 10 cells buffer ($320\,\mathrm{m})$). Table 1 shows a slight improvement in the third significant digit, but Figure 9 shows that the improvement is not consistent across all time segments.
The results indicate some sensitivity to the choice of flexibility in the model, and less to the size of the buffer region. Sensitivity to flexibility is not unexpected. Fuglstad et al., 2015b has previously discussed that more flexibility is not always better, and that a key point is to put the extra flexibility in the correct place.
@@ 6 Discussion
Stationary models are contained in the new model class by using one parameter per coefficient instead of basis functions. The initial state of the spatio-temporal processes can be visualized through the marginal variances and correlations in the initial distribution for the GRF, and the dynamics can be understood through the dampening, the diffusion tensor, and the advection vector field. This means that we can visualize the spatio-temporal dynamics by visualizing a spatially varying dampening, spatially varying ellipses showing the local strength of diffusion in each direction, and a vector field visualising the directions and strengths of advection. In this way, stochastic behaviour can be understood through physical phenomena such as diffusion and advection. Further, in the advection-dominated simulation study, we do not observe the instability seen by Clarotto et al., (2024) for their FEM approximation. This suggests that FVM is a promising alternative to FEM for this problem.
@@
Optimizing our model presents several challenges and opportunities for improvement. One notable limitation is that the vector field describing the anisotropic diffusion exhibits two equivalent optima, complicating optimization efforts, particularly since this vector field varies spatially. Exploring alternative parameterizations of the vector field that circumvent this issue such as Llamazares-Elias et al., (2024) could be beneficial. In our efforts to optimize the model parameters, we have utilized the ADAM optimizer (Kingma and Ba,, 2017), chosen for its effectiveness in handling vanishing gradients—a common issue in machine learning scenarios where certain parameter gradients are extremely small. We have tested several optimizers, including Rmsprop, Adagrad, and Adadelta (Ruder,, 2017), but none outperformed ADAM in our studies. We also experimented with AdaHessian (Yao et al.,, 2021), which scales gradients using the diagonal of the Hessian matrix. However, this approach did not enhance the optimization process, likely because the parameters are too interdependent to benefit from scaling by only the diagonal elements. Calculating the full Hessian matrix, while potentially more effective, is prohibitively expensive in computational terms for this model. Exploring these optimization challenges further could yield significant advancements in the performance and applicability of our approach.
@@
Appendix A Discretizing the advection-diffusion SPDE
@@ A.1 Problem formulation
$\left(-\mathbf{H}(\bm{s})\nabla u(\bm{s},t)+\bm{\omega}(\bm{s})u(\bm{s},t)\right)\cdot{\bm{n}}(\bm{s})=0,\quad\bm{s}\in\partial\mathcal{D},\quad t\in\mathcal{T},$
@@
We solve the advection-diffusion-reaction SPDE given in Equation (S1) numerically by first discretizing in time by backward Euler, and then in space using a finite volume method. This is inspired by Clarotto et al., (2024), but differ in that they used a FEM discretization in space.
@@ A.2 Temporal discretization
$\phantom{=}-\Delta t\left\{\left[\kappa(\cdot)^{2}-\nabla\cdot\mathbf{H}(\cdot)\nabla\right]u^{n+1}(\cdot)+\nabla\cdot(\bm{\omega}(\cdot)u^{n+1}(\cdot))\right\}+$